# Richard A Berk

### v2

@@ Abstract
This paper provides forecasts of rare and extreme heat waves coupled with estimates of uncertainty. Both rest in part on gaining a better understanding of the similarities and differences between hot days under normal circumstances and rare, extreme heat waves. We analyze AIRS data from the American Pacific Northwest and AIRS data from the Phoenix, Arizona region. Forecasting accuracy is excellent for the Pacific Northwest and is replicated for Phoenix. Uncertainty is estimated with nested conformal prediction sets, provably valid even for finite samples. The prediction sets are precise, and the coverage probabilities high. Complications from endogenous sampling are discussed.
Keywords: heat waves, forecasting, AIRS data, machine learning, genetic algorithms, endogenous sampling
@@ 1 Introduction
Extreme heat waves are relatively rare events. But of late, extreme heat waves seem to be increasing in frequency and intensity (Mann et al., 2018). Schmidt (2024) argues that “heat anomalies” may mean that climate science has underestimated the speed at which disastrous global warming will materialize. Many of the consequences for public health and local ecosystems are already troubling (Barriopedro et al., 2023). Some projected consequences are dire (Khatana et al., 2024). The mass media also has taken note. From the Economist, for example, “Deadly heat is increasingly the norm, not an exception to it.” (Economist, June 6, 2024).
@@
Existing studies seeking to understand extreme heat waves have so far produced mixed results (Mann et al., 2018; McKinnon and Simpson, 2022, Fischer et al., 2023; Zhang et al., 2024; Li et al., 2024). Part of the problem is that during the warm-weather months, sequences of hot days have long been “normal.” These spells have some of the features of heat waves that some have linked to climate change. But they arguably are not the rare and extreme heat waves about which there has been so much recent concern.
@@
In this manuscript, the primary goal is to accurately forecast rare and extreme heat waves. To that end, we use novel methods in an effort to help distinguish between usual summer days and the rare, extreme heat waves, perhaps induced by climate change. The data used are a Level 3 product of the NASA Atmospheric Infrared Sounder (AIRS) instrument aboard the Aqua spacecraft (Aumann et al., 2003; Pagano et al., 2010). Observations from the June 2021 heat wave in the North American, Pacific Northwest (PNW) provide an initial forecasting target. Observations from the 2023 July heat wave in the greater Phoenix, Arizona area allow for a possible replication. For both sites, comparisons constructed from “usual” summer days are taken from other months or other years for which no credible claims of extreme temperature episodes have been made. We achieve uncertainty guarantees using nested conformal prediction sets (Kuchibhotla and Berk, 2023).
@@
Section 2 addresses the methodological challenges raised by efforts to forecast extreme heat waves, typically characterized by temperatures that likely fall in the far right tail of temperature distributions; they represent rare events that commonly cause serious data sparsity problems. Section 3 provides a grounded discussion of the remotely sensed, AIRS grid cell data to be used and in Section 4, the research designs that can work well with rare events like extreme heat waves are introduced. Substantial progress is made on data sparsity and on design-based adjustments to mitigate the impact of nuisance variables. Section 5 uses gain scores to examine the marginal, ground-level temperature distributions of an actual heat wave compared to a faux heat wave. The distributions are rather different. Section 6 builds on several variants of crossover designs, with grid cells as the study units, to determine how well machine learning can fit ground-level temperatures during an extreme heat wave compared to ground-level temperatures from ordinary hot summer days. We find that machine learning performance is good, and there is some evidence that the dominant predictors of heat wave temperatures differ somewhat from the dominant predictors of “normal” hot summer temperatures. Section 7 introduces a genetic algorithm to construct a synthetic population of ideal-type, extreme heat wave grid cells, and a synthetic population of ideal-type, ordinary high temperature summer grid cells. Comparisons between the predictor values in these different populations enable a simple form of feature engineering resulting in variable selection. Section 8 uses the selected predictors to forecast a binary response variable denoting whether or not a grid cell experienced an extreme heat wave. Forecasting skill is very good, and nested conformal prediction sets provide provably valid estimates of uncertainty. Section 9 reports a replication of the Pacific Northwest results using an extreme heat wave in the greater Phoenix area. As one would expect, the findings differ a bit, but the overall comparability of the broad conclusions is promising. Section 10 returns to the structure of our research designs in the context of endogenous sampling. We consider the degree to which our efforts to overcome the rare event problem have introduced difficulties in the reported results. We suggest that there is promise in a reformulation from minimizing loss to minimizing risk. Section 11 offers a brief summary of the results and some overall conclusions.
2 Methodological Challenges
@@
In the pages ahead, we use the term “heat wave” to denote extreme heat waves that are relatively rare and arguably enabled by climate change. We use the term “heat dome” as well to capture some instructive climate science thinking (e.g., Jain at el., 2024). Within this framing, there can be many definitions that differ at least in important details (Perkins, 2015). The definitions we use are provisional, constrained in part by available information in the AIRS data.
@@
Almost regardless of how heat waves are defined, a major statistical challenge is data sparsity because extreme heat waves are relatively unusual. One consequence is that very accurate but trivial forecasts can be guaranteed. If in most remotely sensed grid cells most of the time, there is no heat wave, a forecast of no heat wave will almost always be correct without needing to bother with predictors. When predictors are available, they typically will fail to improve the fit or forecasts because both are nearly perfect already. A more subtle consequence is that the design space defined by the predictors will be empty in many regions, making it nearly impossible to estimate well many kinds of nonlinear functions. The fallback of linear statistical methods can be overmatched and miss very important heat wave characteristics. Some consequences of these and related difficulties are addressed later.
@@
Current global climate models (GCMs) do not appear to offer much help. For the spatial scales at which global climate models function, heat waves are generally below the spatial scales that can be well resolved. Various forms of downscaling seem necessary to construct regional climate models (RCMs), but these introduce a number of new challenges (Gebrechorkos et al., 2023; Lafferty and Sriver, 2023; Schneider et al., 2023). As Gettelman and Rood (2016) write, “For these [smaller scale] processes, we often must use statistical treatments to match observations to functions that can be used to describe the behavior. Some processes can be represented by basic laws, other processes must (at the scale of a climate model) be represented by statistical relationships that are only as good as our observations.” To this we add: and only as good as the statistical models themselves (Breiman, 2001b). Moreover, capitalizing on “basic laws” depends on knowing the relevant physics in particular, complex settings, getting boundary conditions from a GCM that are approximately correct, and having access to significant computational resources. And even then, the spatial scale may be too coarse. The analyses to follow suggest that certain quasi-experimental research designs coupled with that machine learning algorithms can help.
@@
There are also difficulties in how best to represent uncertainty. Because there is apparently no consensus model for extreme heat waves, there is no widely acceptable way to formulate a generative model for how the observations materialize. Without such a formulation, the usual uncertainty calculations risk being little more than hand waving. Gettelman and Rood (2016) explain that in climate science, uncertainty is caused by (1) model uncertainty, (2) scenario uncertainty and (3) initial condition uncertainty. In statistics and econometrics, the first is called “model misspecification” (Freedman, 2009). The second is closely related to various forms of stationarity in statistical time series analysis (Box et al., 2016). The third overlaps in statistics with the literature on errors in variables (Fuller, 1987). How these and other statistical and econometric formulations map to those in climate science is largely unresolved.
@@
Further confusion sometimes is introduced by different meanings of the term “model” (Berk et al., 2024). In the natural sciences, a model typically is intended to facilitate explanation and, often, causal inference. The focus is on the manner in which some empirical phenomenon works, and understanding is commonly the dominant goal. Such models necessarily introduce simplifications to further understanding. There also are various other ways models systematically can be right or wrong.
@@
In addition, there can be confusion when an algorithm is called a “model.” “An algorithm is nothing more than a very precisely specified series of instructions for performing some concrete task” (Kearns and Roth, 2019). The goal is to perform that task well without necessarily representing the underlying subject-matter mechanisms. Algorithms are neither right nor wrong; as a technical matter, there is no such thing as a misspecified algorithm. Either an algorithm performs in a satisfactory fashion or it does not. For example, a pattern recognition algorithm might be used to identify malignancies in lung X-rays and often can perform better at that task than a radiologist (Li et al., 2022). But, there is in the algorithm no necessary explanatory or causal content for how the malignant tissue originated and grew.
Frequentist statisticians often use the term “model” in yet another way (Berk et al., 2024). A model usually conveys how the data were generated. Uncertainty commonly begins with random variables and, therefore, the means by which the data were realized. Sometimes the data are realized through design-based approaches, as in various forms of probability sampling used estimate parameters of populations or in random assignment used to estimate causal effects in experiments. Then, a climate science narrative usually is a subject-matter overlay. Sometimes the data are assumed to be realized through processes characterized by subject-matter models as just described. Such models often build on climate science theory. Climate science examples include applications of the generalized extreme value distribution and all manner of linear regression, sometimes hand tailored for time series. All of the usual concerns about model misspecification commonly arise (Berk et al., 2019). In the pages ahead, we will try to make clear when using models, which kind of model we have in mind.
A more subtle potential problem with all models is that as a formal matter, a model should be fully determined before the data are examined. Yet, much model specification is based on model foraging, often as recommended practice. Widely cited treatments of the generalized extreme value distribution (Coles, 2001; Gilleland and Katz, 2016), for instance, provide a variety of statistical tools for empirical determination of which special case of the umbrella formulation that should be employed: the Gumbel, Fréchet or Weibull distribution (Kotz et al., 2000). The result is sanctioned cherry picking, which can invalidate all statistical inference (Leeb and Pötscher, 2005; 2006; 2008).
Post-model-selection challenges follow that if acknowledged, often can be usefully addressed (Berk et al., 2013). We illustrate a form of post-model selection inference (PoSI) later using conformal prediction sets when we deliver provably valid uncertainty assessments applying nested conformal prediction sets to data restructured within a crossover design. Consistent with recent developments in statistics, design-based uncertainty estimates often accomplish what model-based uncertainty assessments cannot (Breiman, 2001b; Rubin, 2008; Rinaldo et al., 2019). We try to capitalize on design-based approaches in the analyses to follow.
3 AIRS Remotely Sensed Data
We provide next a relatively deep dive into the production of remotely sensed data from the Aqua spacecraft. Perhaps the primary message is that complicated hardware and software are required on which researchers must rely. There are at this point no “do-overs.” Moreover, the “Level 3” data fundamental to our analyses has the smallest spatial scale at the requisite geographical locations for remotely sensed data readily available in the public domain. Table 1 shows the variable downloaded for the analyses to follow.
The Aqua spacecraft responsible for the AIRS data was launched into polar orbit at an altitude of 705 km on May 4, 2002. Acquiring and processing AIRS data began in September of that year. Aqua crosses the equator during the ascending, or the northward daytime part of its orbit, at 1:30 pm local time. The descending, or southward, nighttime, part of its orbit crosses the equator at 1:30 am local time. Each half-orbit takes roughly 45 minutes, so local time of data acquisition is always within 22.5 minutes of equator crossing. AIRS successively scans across its 1500 km field of view taking data in 90 circular or elliptical footprints. As the spacecraft advances, the sensor resets and obtains another scan line. 135 scans are completed in six minutes, and this 90 by 135 footprint spatial array constitutes a “granule” of data. AIRS collects 240 granules per day, 120 on the daytime, ascending portions of orbits and 120 on nighttime, descending portions.
At each of the 12,150 footprints in a granule, AIRS observes Earth and its atmosphere in 2378 infrared spectral channels. Roughly speaking, the channels sense the surface and different altitudes in the atmosphere. The instrument counts photons at the different wave numbers (inverse wavelengths). These counts are converted to brightness temperatures ranging from zero to about 340 degrees Kelvin. Such data are called Level 1 products (Aumann et al., 2020). Certain atmospheric characteristics are related to photon emission, and these characteristics can be retrieved by solving complex sets of equations. The computational implementation is called a retrieval algorithm (Susskind et al., 2020), which converts spectra of brightness temperatures into estimates of geophysical variables. For us, the potentially relevant variables include surface temperature, tropopause height, and vertical profiles of air temperatures and water vapor mass-mixing ratios at various altitudes starting from the near-surface up to 250 mb pressure. These data are known as Level 2 products (Thrastarson, 2021).
The mission also provides a globally gridded version, summary product that averages Level 2 data on a one-degree-by-one degree daily spatial grid. These are Level 3 data products (Tian et al., 2020) intended for users who do not require the higher spatial and temporal resolution of Level 2. Typically studies of global climate processes use Level 3 because it is easier to compare with the output of climate models. Many variables are provided in the AIRS Level 3 daily product. The ones of interest to us for this study are shown in Table 1. Their variable names are “official” in that those are the names that come with the downloaded data. We shorten some names and change some names to facilitate data processing and the construction of tables and figures. But the mapping from the names in Table 1 to names used in subsequent tables and figures should be clear.111 In Table 1, the phrase “Pressure levels of temperature and water vapor profiles” means that atmospheric pressure is proxy measure for altitude for both temperature and the water vapor profile. The former is in degrees Kelvin. The latter is a lot like relative humidity at different altitudes but is the proportion of the total atmospheric profile that is water vapor at a given altitude. The total atmospheric profile is the proportional representation of different constituents of the atmosphere at a given altitude such as oxygen, carbon dioxide, nitrogen, and water vapor.
Name Dimensions Description
Latitude $360\times 180$ Latitude at the center if the grid cell
Longitude $360\times 180$ Longitude at the center of the grid cell
LandSeaMask $360\times 180$ 1 = land, 0 = sea
Topography $360\times 180$ Topography of the Earth in meters above the geoid
StdPressureLev $24\times 1$ Pressure levels of temperature and water vapor profiles
(we use only 1 – 12). The array order is from the surface
@@ 3 AIRS Remotely Sensed Data
SurfAirTemp $360\times 180$ Temperature of the atmosphere at the Earth’s surface in
@@
TropHeight $360\times 180$ Height of the tropopause in meters above sea level
Temperature $360\times 180\times 24$ Temperature of the atmosphere in degrees Kelvin (we
use only bottom 12)
H2O MMR $360\times 180\times 12$ Water vapor mass-mixing ratio in grams/kilogram
Table 1: AIRS Level 3 variables used in this study, along with their dimensions and brief descriptions.
@@ 4 Research Design
The research design is influenced substantially by the rare events problem. We capitalize on crossover, quasi-experimental principles. The study units serve as their own controls. Forecasting skill likely is the beneficiary. They also provide a foundation for overcoming statistical problems from rare events.
To begin, any research design in service of forecasts of extreme heat waves must address the question “compared to what?” The obvious answer is compared to “non-heat waves.” But the details can be subtle. Summer months in many parts of the world can manifest several very warm days in a row that may be unusual, and when juxtaposed to recent history, may seem extreme.
@@
We begin with the 2021 Pacific Northwest extreme heat wave affecting parts of the U.S. and Canada. During the last four days of June, the Pacific Northwest experienced what many observers called an extreme heat wave. Although there are debates over what exactly constitutes an extreme heat wave, this event seems to qualify from almost any informed perspective. With so many extant definitions (Perkins, 2015, Barriopedro et al., 2023), and nothing close to a consensus on best practices, we will provisionally operationalize an extreme heat wave as a single discrete high temperature event over several days under a slowly moving “heat dome” (Li et al., 2024). Other formulations certainly are possible, and perhaps even preferable, depending on the policy and/or research questions asked.222 The many possible definitions of a heat wave leads to some clumsy wording. The term “true” heat wave implies broad definitional agreement. The same applies to “actual” heat wave. The use of “extreme” heat wave can be quite vague unless “extreme” is defined. Even the term “heat wave” also can mean that we are using a widely accepted definition. We use all four at various times depending on context with the understanding that the definition just provided is what we intend to convey.
@@
Figure 1: U.S. Pacific Northwest study region shown by the large red rectangle. Smaller rectangles are grid cells.
@@
Media coverage of the June heat wave was by and large consistent with data from the Community Earth System Model (https:// www.cesm.ucar.edu/), which is not employed in the current analysis. All relevant June days were confirmed with the PNW AIRS data used in the analyses to follow. The same sources were used to identify the absence of extreme heat waves in manner described shortly.333 The spatial boundaries of an extreme heat wave can be difficult to determine. One complication is a heat dome responsible can cover many square kilometers and move slowly over time. Different fractions of potentially affected grid cells, nominally in the same locale, will fall under the dome from day to day. In the case of the PNW June 2021 heat wave, the four-day definition seems to be a reasonable approximation.
@@
Figure 1 shows the location of our study area: a $12^{\circ}\times 12^{\circ}$ region over the northwest U.S. and southeastern Canada. For this analysis, we use daily data from these 144 grid cells, each grid cell indexed by location and day. The widely varying topography apparent from Figure 1 can be associated with very different microclimates that must be taken into account as we proceed. Clearly, surface temperatures can be quite dissimilar spatially whether or not there is an extreme heat wave. We will later consider similar data from the greater Phoenix area.
Figure 2: The extreme heat wave is compared separately to Faux Heat Waves 1 and 2. The top blue line shows the research design for the extreme heat wave. The middle blue line shows the research design Faux Heat Wave 1. The bottom blue line shows the research design for Faux Heat Wave 2. Each design has a pre-test, post-test, and predictors. The extreme heat wave is for June, 2021. Faux Heat Wave 1 is shifted back 4 Days, and Faux Heat Wave 2 is shifted ahead 30 days. The pre-tests and post-tests are each computed separately over grid cells as the surface temperature mean of 4 consecutive days.
@@
To what should the PNW 2021 heat wave been compared? Our research design for the PNW AIRS data is shown in Figure 2: the top blue line for the extreme heat wave, the middle blue line for comparison “Faux Heat Wave 1,” and the bottom blue line for comparison “Faux Heat Wave 2.” The study units are the same 144 grid cells for the extreme heat wave and both faux heat waves. Much of the data for a grid cell can differ depending on the day on which remote sense data were collected. (See Table 1.)
Starting with the heat wave at the top representing the month of June, 2021, the post-test is the mean surface temperature for each grid cell over the 4 last days of the month. The pre-test is the mean surface temperature for each grid cell over the 4 sequential days immediately preceding the first post-test day. In both cases, means are used to help remove random variation in noisy measures of daily surface temperatures. Having a pre-test as well as a post-test confers benefits discussed in the next section. All of the predictor values are taken from 14 days before the beginning of the true heat wave. The same 144 grid cells constitute the post-test, pre-test, and lagged predictors.
@@
Faux Heat Wave 1, shown as the middle blue line, has the same design configuration as the heat wave, but is shifted back in time 4 days before the extreme heat began. Its post-test does not include the extreme heat wave, which provides one answer to the question “compared to what?” Faux Heat Wave 2 at the bottom of Figure 2 also has the same design configuration as the extreme heat wave, but is shifted forward in time one month to July, 2021 such that the extreme June heat wave also is not included in the post-test. It provides as second answer to the question “compared to what?” For both faux heat waves, the same 144 grid cells constitute the post-test, pre-test, and lagged predictors.
@@
Faux Heat Wave 1 has the advantage of a post-test falling immediately before the extreme heat wave post-test so that potential weather confounders might be quite similar. The risk is that some of the extreme heat wave characteristics may be in part a spillover from the immediately preceding several days; a larger temporal separation might be desirable. Therefore, Faux Heat Wave 2 provides a temporal separation of a month, but risks that summers in July may on the average differ from summers in June in ways confounded with surface temperatures.444 We also considered computing the pre-test using the grid cell means for the full set of June days before the heat wave. However, this approach builds in temporal warming trends in surface temperatures gradually unfolding during the June.
5 Surface Temperature Changes from Pre-Test to Post-Test

### v3

@@ Abstract
This paper provides some foundations for valid forecasting of rare and extreme heat waves through a better understanding of the similarities and differences between several consecutive hot days under normal circumstances and rare, extreme heat waves. We analyze AIRS data from the American Pacific Northwest and AIRS data from the Phoenix, Arizona region. A genetic algorithm is used to help determine the most promising predictors. Classification accuracy with supervised learning is excellent for the Pacific Northwest and is replicated for Phoenix. Conformal prediction sets are considered as a way to represent forecasting uncertainty. Complications caused by endogenous sampling are discussed.
keywords:
heat waves, cross-over design, forecasting, AIRS data, supervised machine learning, genetic algorithms, spatial dependence, endogenous sampling
@@ 1 Introduction
Extreme heat waves are relatively rare events. But of late, extreme heat waves seem to be increasing in frequency and intensity (Mann et al., 2018). Despite a growing literature (Marx et al., 2021), Schmidt (2024) argues that “heat anomalies” may mean that climate science has underestimated the speed at which disastrous global warming will materialize. Many of the consequences for public health and local ecosystems are already troubling (Barriopedro et al., 2023). Some projected consequences are dire (Khatana et al., 2024). The mass media also has taken note. From the Economist, for example, “Deadly heat is increasingly the norm, not an exception to it.” (Economist, June 6, 2024).
@@
Existing studies seeking to understand extreme heat waves have so far produced mixed results (Mann et al., 2018; McKinnon and Simpson, 2022, Fischer et al., 2023; Zhang et al., 2024; Li et al., 2024). Part of the problem is that during the warm-weather months, sequences of hot days have long been “normal.” These spells have certain features of heat waves that some have linked to climate change (Bartusek, 2022; Dosio et al., 2028). But they arguably are not the rare and extreme heat waves about which there has been so much recent concern.
@@
In this manuscript, the primary goal is help built a foundation on which valid forecasts of rare and extreme heat waves can be made. To that end, we use a within-subject, quasi-experimental design combined with algorithmic methods in an effort to better distinguish between ordinary, hot summer days and rare, extreme heat waves, perhaps induced by climate change. The data used are a Level 3 product of the NASA Atmospheric Infrared Sounder (AIRS) instrument aboard the Aqua spacecraft (Aumann et al., 2003; Pagano et al., 2010). Observations from the June 2021 heat wave in the North American, Pacific Northwest (PNW) provide an initial forecasting target. Observations from the 2023 July heat wave in the greater Phoenix, Arizona area allow for a replication. For both sites, “control group” comparisons constructed from “usual” summer days are taken from other months or other years for which no credible claims of extreme temperature episodes have been made.
@@
Section 2 addresses some methodological challenges raised by efforts to forecast extreme heat waves, typically characterized by temperatures that likely fall in the far right tail of temperature distributions; they represent rare events that commonly cause serious data sparsity problems. Section 3 provides a grounded discussion of the remotely sensed, AIRS grid cell data to be used and in Section 4, the research designs that can work well with rare events like extreme heat waves are introduced. Substantial progress is made on data sparsity and on design-based adjustments to mitigate the impact of nuisance variables. Section 5 uses gain scores to examine the marginal, ground-level temperature distributions of an actual heat wave compared to a faux heat wave. The distributions are rather different. Section 6 examines whether the predictor distribution differ as well. They do. Section 7 introduces a genetic algorithm to construct a synthetic population of ideal-type, extreme heat wave grid cells, and a synthetic population of ideal-type, ordinary high temperature summer grid cells. Comparisons between the predictor values in these different populations enable a simple form of feature engineering for variable selection. Section 8 uses the selected predictors to examine the promise of forecasting a binary response variable denoting whether or not a grid cell experiences an extreme heat wave. Section 9 reports an effort to reproduce the Pacific Northwest results using an extreme heat wave in the greater Phoenix area. As one would expect, the results somewhat differ, but the overall comparability of the broad conclusions is encouraging. Section 10 returns to the structure of our research designs and the matter of endogenous sampling. We consider the degree to which our efforts to overcome the rare event problem have introduced difficulties in the reported results. We suggest that there is promise in a reformulation from minimizing loss to minimizing risk, but subject-matter and policy considerations are fundamental. Section 11 offers a brief summary of the results and some overall conclusions. The findings are framed as highly provisional in part because many more heat waves in a variety of locations need to be studied.
2 Some Methodological Challenges
@@
In the pages ahead, we use the term “heat wave” to denote extreme heat waves that are relatively rare and arguably enabled by climate change. We use the term “heat dome” as well to capture some instructive climate science thinking (e.g., Jain at el., 2024). Within this framing, there can be many definitions that differ at least in important details (Perkins, 2015; Piticar et al., 2019). The definitions we use are provisional, constrained in part by available information in the AIRS data.
@@
Almost regardless of how heat waves are defined, a major statistical challenge is data sparsity because extreme heat waves are relatively unusual. One consequence is that very accurate but trivial forecasts can be guaranteed. If in most remotely sensed grid cells most of the time, there is no heat wave, a forecast of no heat wave will almost always be correct without needing to bother with predictors. When predictors are available, they typically will fail to improve the fit or forecasts because both are nearly perfect already. A more subtle consequence is that the design space defined by the predictors will be empty in many regions, making it nearly impossible to estimate well many kinds of nonlinear functions. In other words, for many combinations of predictor values, there are no data. The fallback of linear statistical methods can be overmatched and miss very important heat wave characteristics. Some consequences of these and related difficulties are addressed later.
@@
Current global climate models (GCMs) do not appear to offer much help. For the spatial scales at which global climate models function, heat waves are generally below the spatial scales that can be well resolved. Various forms of downscaling seem necessary to construct regional climate models (RCMs), but these introduce a number of new challenges (Gebrechorkos et al., 2023; Lafferty and Sriver, 2023; Schneider et al., 2023). As Gettelman and Rood (2016) write, “For these smaller scale processes, we often must use statistical treatments to match observations to functions that can be used to describe the behavior. Some processes can be represented by basic laws [of physics], other processes must (at the scale of a climate model) be represented by statistical relationships that are only as good as our observations.” To this we add: and only as good as the statistical models themselves (Breiman, 2001b; Freedman, 2009). Moreover, capitalizing on “basic laws” depends on knowing the relevant physics in particular, complex settings, getting boundary conditions from a GCM that are approximately correct, and having access to significant computational resources. And even then, the spatial scale may be too coarse. The analyses to follow suggest that certain quasi-experimental research designs can help when used together with supervised learning.
@@
There are also difficulties in how best to represent uncertainty. Because there is apparently no consensus model for extreme heat waves, there is no widely acceptable way to formulate a generative model for how the observations materialize. Without such a formulation, the usual uncertainty calculations risk being little more than hand waving. Gettelman and Rood (2016) explain that in climate science, uncertainty is caused by (1) model uncertainty, (2) scenario uncertainty and (3) initial condition uncertainty. In statistics and econometrics, the first is called “model misspecification” (Freedman, 2009). The second is closely related to various forms of stationarity in statistical time series analysis (Box et al., 2016). The third overlaps in statistics with the literature on errors in variables (Fuller, 1987). How these and other statistical and econometric formulations map to those in climate science is largely unresolved. Spatial dependence complicates matters further.
@@
Confusion also can be introduced by different meanings of the term “model” (Berk et al., 2024). In the natural sciences, a model typically is intended to facilitate explanation and, often, causal inference. The focus is on the manner in which some empirical phenomenon works, and understanding is commonly the dominant goal. Such models necessarily introduce simplifications to further understanding. There also are various other ways models systematically can be right or wrong.
@@
An algorithm is not a model. “An algorithm is nothing more than a very precisely specified series of instructions for performing some concrete task” (Kearns and Roth, 2019). The goal is to perform that task well without necessarily representing the underlying subject-matter mechanisms. Algorithms are neither right nor wrong; as a technical matter, there is no such thing as a misspecified algorithm. Either an algorithm performs in a satisfactory fashion or it does not. For example, a pattern recognition algorithm might be used to identify malignancies in lung X-rays and often can perform better at that task than a radiologist (Li et al., 2022). But, there is in the algorithm no necessary explanatory or causal content for how the malignant tissue originated and grew.111 Recent developments in physics-informed algorithms muddies the differences between models and algorithms (Kerniadakis et al., 2021). Formulations from physics, such as certain physical invariants, are embedded in special neural network architectures. Ideally, fitting performance is improved.
3 The AIRS Remotely Sensed Data
The Aqua spacecraft responsible for the AIRS data was launched into polar orbit at an altitude of 705 km on May 4, 2002. Acquiring and processing AIRS data began in September of that year. AIRS observes Earth and its atmosphere in 2378 infrared spectral channels. The initial spectral input is processed and stored as Level 1 products. (Aumann et al., 2020). Level 2 products result when the Level 1 Products are converted into geophyscial variables such as atmospheric moisture content at various altitudes (Suskind et al., 2020; Thrastarson, 2021). Level 3 data products (Tian et al., 2020) are intended for users who do not require the higher spatial and temporal resolution of Level 2. In the analyses to follow, we use Level 3 data.
Name Description
$Latitude$ Latitude at the center if the grid cell.
$Longitude$ Longitude at the center of the grid cell.
$LandSeaMask$ 1 = land, 0 = sea.
$Topography$ Topography of the Earth in meters above the geoid.
$StdPressureLev$ Pressure levels of temperature and water vapor profiles
(we use only 1 – 12). The array order is from the surface.
@@ 3 The AIRS Remotely Sensed Data
$SurfAirTemp$ Temperature of the atmosphere at the Earth’s surface in
@@
$TropHeight$ Height of the tropopause in meters above sea level.
$Temperature$ Temperature of the atmosphere in degrees Kelvin (we
use only 1 – 12). The array order is from the surface.
upward.
[…]
Table 1: AIRS Level 3 variables used in this study and brief descriptions.
The available Level 3 variables are shown in Table 1. Their names are “official” in that those are the names that come with the downloaded data. We later shorten some names and change some names to facilitate data processing convenience and the construction of approachable tables and figures. But the mapping from the names in Table 1 to names used in subsequent tables and figures should be clear.222 In , the phrase “Pressure levels of temperature and water vapor profiles” means that atmospheric pressure is proxy measure for altitude for both temperature and the water vapor profile. The former is in degrees Kelvin. The latter is a lot like relative humidity at different altitudes but is the proportion of the total atmospheric profile that is water vapor at a given altitude. The total atmospheric profile is the proportional representation of different constituents of the atmosphere at a given altitude such as oxygen, carbon dioxide, nitrogen, and water vapor.
[…]
@@ 4 Research Design
Any research design meant to anticipate extreme heat waves must address the question “compared to what?” The obvious answer is compared to “non-heat waves.” We begin with the 2021 Pacific Northwest extreme heat wave that affected parts of the U.S. and Canada. During the last four days of June, the Pacific Northwest experienced what many observers called an extreme heat wave. Although there are debates over what exactly constitutes an extreme heat wave, this event seems to qualify from almost any informed perspective. With so many extant definitions (Perkins, 2015, Barriopedro et al., 2023), and nothing close to a consensus on best practices, we will provisionally operationalize an extreme heat wave as a single discrete, unusually high temperature event over several days under a slowly moving “heat dome” (Li et al., 2024). More details are provided shortly. Other formulations certainly are possible, and perhaps even preferable, depending on the policy and/or research questions asked.333 The many possible definitions of a heat wave leads to some clumsy wording. The term “true” heat wave implies broad definitional agreement. The same applies to “actual” heat wave. The use of “extreme” heat wave can be quite vague unless “extreme” is defined. Even the term “heat wave” also can mean that we are using a widely accepted definition. To reduce stale repetition, we use all four terms at various times depending on context with the understanding that the definition just provided is what we intend to convey.
@@
Figure 1: U.S. Pacific Northwest study region shown by the large red rectangle. Smaller squares are grid cells.
@@
Media coverage of the June heat wave by and large was consistent with data from the Community Earth System Model (https://www.cesm.ucar.edu/), which is not employed in the current analysis. All relevant June days were confirmed with the PNW AIRS data used in the analyses to follow. The same sources were used to identify the absence of extreme heat waves in manner described shortly.
Yet, the spatial boundaries of an extreme heat wave can be difficult to determine. One complication is a heat dome responsible can cover many square kilometers and move slowly over time. Different fractions of potentially affected grid cells, nominally in the same locale, will fall under the dome from day to day. In the case of the PNW June 2021 heat wave, the four-day definition seems to be a reasonable approximation.
@@
Figure 1 shows the location of our study area: a $12^{\circ}\times 12^{\circ}$ region over the northwest U.S. and southwestern Canada. For our analysis, we use daily data from these 144 grid cells, each grid cell indexed by location and day. The widely varying topography is apparent from Figure 1.
@@
To what should the PNW 2021 heat wave been compared? The PNW AIRS curated data structure is shown in Figure 2: the top blue line for the extreme heat wave, the middle blue line for comparison “Faux Heat Wave 1,” and the bottom blue line for comparison “Faux Heat Wave 2.” The study units are the same 144 grid cells for the extreme heat wave and both faux heat waves; the same 144 grid cells are included within each of these settings. For variables not treated as fixed, the data values associated with a grid cell can differ depending on the day on which remotely sensed data were collected.
Figure 2: The extreme heat wave data format at the top compared separately to Faux Heat Waves 1 or 2.
Starting with the heat wave at the top representing the month of June, 2021, the post-test is the mean surface temperature for each grid cell over the 4 last days of the month. The pre-test is the mean surface temperature for each grid cell over the 4 sequential days immediately preceding the first post-test day. In both cases, means are used to help remove random variation in noisy measures of daily surface temperatures or variation too transient to be associated systematically with an extreme heat wave. The terms “pre-test” and “post-test” are common in experiments and observational studies seeking to estimate causal relationships. Our goal is to improve forecasting validity, which does not require estimates of cause and effect. Substantial associations between outcomes and outcome precursors can be sufficient. However, causal reasoning can often help to better formulate how the data are curated and organized.
Having a pre-test as well as a post-test confers benefits discussed in the next section. Based on past research (Li et al., 2034), all of the predictor values are taken from 14 days before the beginning of the true heat wave. We consider whether they are promising extreme heat wave precursors 2 weeks in advance. Other lag sizes are certainly possible and worth exploring. The same 144 grid cells constitute the post-test, pre-test, and lagged predictor observations.
@@
Faux Heat Wave 1, shown as the middle blue line, has the same design configuration as the actual heat wave, but is shifted back in time 4 days before the extreme heat began. Its post-test does not include the extreme heat wave, which provides one answer to the question “compared to what?” Faux Heat Wave 2 at the bottom also has the same design configuration as the extreme heat wave, but is shifted forward in time one month to July, 2021, such that the extreme June heat wave also is not included in the post-test. It provides as second answer to the question “compared to what?” For both faux heat waves, the same 144 grid cells constitute the post-test, pre-test, and lagged predictor observations.
@@
Faux Heat Wave 1 has the advantage of a pre-test falling immediately before the extreme heat wave post-test so that potential confounders might have quite similar values. The risk is that some of the extreme heat wave characteristics may be in part a carried over from the immediately preceding several days; a larger temporal separation might be desirable (Benjamin-Chung et al., 2018). Therefore, Faux Heat Wave 2 provides a temporal separation of a month, but risks that summers in July may on the average differ from summers in June in ways confounded with surface temperatures.
In summary, the data generation thought experiment can be operationalized as Figure 2. Figure 2 can be implemented using certain features of within-subject quasi-experimental designs (Maxwell et al., 2018). In particular, study units can be used as their own controls. We begin to capitalize on this in the next section.
5 Temperature Changes from Pre-Test to Post-Test